If Jen's current age is 32 and Devin's current age is 6. Write an equation which models how old, in years, Jen will be when she is 3 times as old as Devin.
Answer by gwendolyn(128)

First, we'll assign Jen and Devin's ages variables.
let J=Jen's age
let D=Devin's age
Next, we'll write the equations for Jen's and Devin's ages.
J=32
D=6
We need to know how many years it will be until Jen's age is three times Devin's.
We'll assign a variable to this amount of time.
let x=the number of years until Jen's age is 3 times Devin's
Then, we can write the equation that will satisfy the question.
Jen's age will be 3 times that of Devin's in x years, so:
(J+x)=3(D+x)
After that, we can substitute in the values of J and D:
32+x=3(6+x)
Following this, we distribute the 3 over the parentheses
32+x=18+3x
We could use this as the equation, but I will solve it just in case you need me to.
To solve the equation, we subtract 18 from both sides:
14+x=3x
And then we subtract x from both sides:
14=2x
Finally, we divide both sides by two to isolate the variable.
7=x
x=7
Therefore, in seven years, Jen's age will be triple Devin's.